OpenStudy (anonymous):

A particle moves along a straight line with equation of motion s=t^4-1t^3. Find the value of t (other than 0) at which the acceleration is equal to zero

OpenStudy (anonymous):

I get up to 4t^3-3t^2 => 12t^2-6t. then what?

OpenStudy (anonymous):

acceleration = second derivative of s. next =0

OpenStudy (anonymous):

v=s'=4t^3-3t^2 a=12t^2-6t=0 2t^2-t=0 t(2t-1)=0 t=0 or t=1/2

OpenStudy (anonymous):

a=s''(t)

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

yes! thanks :]

OpenStudy (anonymous):

welcome!